Aki's bicycle designs has determined that when x hundred bicycles are built, the average cost per bicycle is given by C(x)=0.2x^2-0.1x+3.799, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the avearge cost per bicycle?
Answer by stanbon(75887)

Aki's bicycle designs has determined that when x hundred bicycles are built, the average cost per bicycle is given by C(x)=0.2x^2-0.1x+3.799, where C(x) is in hundreds of dollars. How many bicycles should the shop build to minimize the average cost per bicycle?
--------------------------
Minimum occurs when x = -b/(2a) = 0.1/(2*0.2) = 1/4
---
# of bikes should be (1/4)*100 = 25
=========================================
Cheers,
Stan H.
===========